# neurips_2025.bib

### v1

@@ 4.3 Ablations Study
Ablations on post-training strategies. In Section 4.3, we evaluate the performance of our multi-frame post-training strategy by comparing the following settings: (1) Basic-Post: Starting from our basic single-frame VLA model and being further post-trained discretely without multi-frame modeling. (2) + Multi-frame: Built on the Basic-Post by directly adding multiple frames as input to the vision-language backbone and post-trained in the discrete action space. (3) + Trans.&Multi-frame: Transferring the discrete action prediction to motion feature prediction, but adding multi-frame inputs to the vision-language backbone instead of the decoder. (4) Ours: Our full pipeline that models multi-frame awareness in the decoder. All settings are post-trained on the same datasets, and the past frame number of (2)~(4) is 6. As shown in Section 4.3, Basic-Post achieves a modest overall success rate of 31.0% with 5.18 Hz. Introducing multi-frame inputs without architectural support, + Multi-frame only provides a marginal improvement (+1.4%), but results in a substantial drop in inference speed (3.09 Hz, a 40.3% decrease), suggesting that naively adding multiple frames may be difficult to obtain an obvious gain even feeding the same input. In contrast, with the transfer from the discrete action pertaining and naive multi-frame modeling, + Trans.&Multi-frame boosts performance to 68.1%, but still slows the inference speed (5.03 Hz). Finally, our full method achieves the best overall performance (70.9%) and considerably improves inference frequency to 8.73 Hz (+68.5%). This shows that our method not only enhances performance but also improves inference efficiency.
Table 3: Ablation on post-training strategies. The average success rates across all 12 tasks in SimplerEnv and the inference speed are shown in this table, based on CronusVLA 7B.
# Methods Overall (%) speed (Hz)
1 Basic-Post 31.0 5.18
2 + Multi-frame 32.4 3.09
3 + Trans.&Multi-frame 68.1 5.03
4 Ours 70.9 8.73
Table 4: Ablation on cross-frame decoder. Based on ours 7B with a past frame number of 6, all designs are evaluated in SimplerEnv, including G-VM, G-VA, and W-VM, following  Section 4.1.
# Method G-VM G-VA W-VM Ave.
@@
2 w/o. cross-atten 76.4 71.2 57.3 68.3
3 flow matching 75.0 69.5 58.4 67.6
@@
Figure 6: Extended ablation studies. (a) Varying the number of input frames affects overall success rate and inference speed. (b) The effect of multi-frame regularization on performance, which is across all SimplerEnv settings and based on CronusVLA 7B. (c) Probability density curves of action similarity, obtained via kernel density estimation on 25% randomly sampled training data of LIBERO.
@@
Ablation on cross-frame decoder. In  Section 4.3, we present more exploration of architectural variations within the cross-frame decoder. The w/o. modulator setting omits the modulator, treating current (1 frame) and past motion features (6 frames) indiscriminately, which substantially degrades performance due to the potential dominance of irrelevant historical information over critical current cues. For w/o. cross-atten, we remove the cross-attention mechanism of our original design and replace it with a direct self-attention network. It shows our design achieves better overall performance and a higher per-task success rate. Notably, self-attention leads to quadratic growth in computation with past frames increasing, whereas our cross-attention approach scales linearly, enabling more efficient support for long-horizon modeling. We also investigate a flow matching approach [flowmatching], substituting the diffusion objective with a flow matching objective, as shown in Section 4.3 #3.
@@
The impact of frame number. In this section, we analyze how varying the number of input frames affects different multi-frame modeling strategies. Given that tasks differ in their reliance on temporal information, we conduct a representative analysis in the SimplerEnv benchmark. The study compares CronusVLA 7B, CronusVLA 0.5B, and a baseline model (Basic-Post) with a naive multi-frame extension. As shown in Figure 6 (a), results show that increasing the total frame count (the current one frame plus past frames) from 2 to 9 yields different gains in average success rate across 2.5k trials on SimplerEnv. More frames do not consistently lead to better outcomes: CronusVLA 7B performs best with 7 frames, while the 0.5B variant performs best with 4, suggesting that a moderate amount of temporal information can enhance performance, whereas excessive temporal input may lead to performance degradation. Compared with Basic-Post, CronusVLA maintains high inference speed across frame counts, avoiding excessive latency overhead.
Ablation on multi-frame regularization. We find that multi-frame regularization facilitates more accurate action learning and improves model convergence. We attribute this to the property that, during post-training, our multi-frame regularization strategy effectively preserves the single-frame perception capability while allowing sufficient training of the cross-frame decoder. As a result, it consistently outperforms the variant without regularization on SimplerEnv, as shown in Figure 6 (b). Additional experiments on how it affects the model convergence are provided in Section B.1.
Table 5: Ablation on action adaptation mechanism. With fixed 10k training steps on CronusVLA 7B, FT indicates further finetuning, while Adapt. denotes whether action adaptation is employed.
FT Adapt. Spatial Object Goal Long Ave.
88.1 90.6 86.5 63.4 82.2
87.6 91.6 86.6 63.6 82.4
90.1 94.7 91.3 68.7 86.2
Ablation on action adaptation. (1) Action similarity. We compare the action similarity between multi-frame visual features extracted by the vision encoder, as well as the action similarity between the modulated features. We retrieve the top-6 most similar candidates across different episodes and analyze the distribution of their corresponding action similarities, as shown in Figure 6 (c). It shows that our modulated features exhibit a more concentrated distribution in the 0.9 ~1.0 similarity range, indicating a higher likelihood of retrieving accurate actions. This suggests that composing motion features across multiple timesteps is more effective for action retrieval than directly using multi-frame visual features, as the latter are more susceptible to interference from irrelevant backgrounds or fine-grained visual details. (2) Action adaptation mechanism. As shown in Section 4.3, after training CronusVLA for a fixed number of steps, the model achieves an average success rate of 82.2% on LIBERO. Incorporating our action adaptation and continuing to train both the action adapter and decoder further improves performance to 86.2%. In contrast, continually finetuning the entire language model and decoder under the same training budget introduces performance variance to 82.4% without consistent improvement. To reduce stochastic noise, we report results averaged over repeated evaluations. The results demonstrate the effectiveness of our action adaptation mechanism. Other ablation studies are included in Appendix B.
@@ 5 Conclusion
Appendix
Appendix A Case Study for Long-horizon Tasks
A.1 Real-world Experiments

### v2

@@ 4.3 Ablations Study
Ablations on post-training strategies. As shown in Tab. 4, we evaluate the effectiveness of different multi-frame post-training strategies (with total 7 frames) by comparing the following configurations: (1) Baseline: Our basic single-frame VLA model, further post-trained without multi-frame modeling. (2) +M.F.: Directly feeding multiple images into the vision-language backbone and post-training in a discrete action space. This yields a margin performance gain (only +4.5%) but considerably reduces inference speed (-40.3%). (3) +M.F. +Dec.: Replacing discrete action prediction with continuous feature prediction, training only the decoder during multi-frame modeling. This setting achieved substantial performance improvement and considerably boosts inference speed (8.73 Hz vs. 3.09 Hz), thanks to the elimination of autoregressive decoding and the caching of feature chunking. (4) +M.F. +Dec. +V.L.: Enabling training of both the decoder and vision-language backbone under multi-frame post-training also improves the overall performance, attributed to the enhanced fitting ability. (5) +M.F. +Dec. +V.L. +Reg. (Ours): Utilizing the multi-frame regularization can effectively leverage single-frame perception of vision-language backbone while enabling sufficient multi-frame modeling of the decoder, leading to superior performance. Additionally, we observe that multi-frame regularization considerably improves the convergence (details in Appendix).
Settings G-VM G-VA W-VM Ave. Speed (Hz)
Baseline 43.3 39.4 10.4 31.0 5.18
+M.F. 45.5 47.4 4.2 32.4 3.09
+M.F. +Dec. 52.2 58.1 34.4 48.2 8.73
+M.F. +Dec. +V.L. 72.7 72.6 56.3 67.2 8.73
+M.F. +Dec. +V.L. +Reg. 78.6 73.8 60.4 70.9 8.73
Table 4: Ablation on post-training strategies. Success rates on SimplerEnv and inference speeds are shown. We discuss adding multiple images, training decoder, training vision-language backbone, and multi-frame regularization.
# Settings G-VM G-VA W-VM Ave.
1 w/o. cross-atten 76.4 71.2 57.3 68.3
@@
3 MLP-based decoder 66.4 53.6 35.4 51.8
4 SiT-based decoder 75.0 69.5 58.4 67.6
@@
Table 5: Ablation on cross-frame decoder designs. Based on CronusVLA 7B with a past frame number of 6.
Figure 5: Impact of frame number. Varying the number of input frames affects the success rate and inference speed.
@@
Ablations on cross-frame decoder designs. As illustrated in Tab.5, we first study the core design in our DiT-based decoder. For w/o. cross-atten, we remove the cross-attention mechanism of our original design and replace it with a direct self-attention network. Our cross-attention design achieves superior overall performance and higher per-task success rates. Notably, self-attention incurs quadratic computational growth with increasing frame numbers, whereas our cross-attention mechanism scales linearly, enabling support for extremely large frame sequences. For w/o. modulator, we omit the modulator, treating current (1 frame) and past learnable features (6 frames) indiscriminately without feature splitting and learnable positional embedding, which substantially degrades performance due to the potential dominance of irrelevant historical information over critical current cues. Then, we explore alternative decoder architectures by adopting an MLP-based decoder with the same parameter size as our DiT-based model. Due to the large-scale post-training data, the limited representational capability of the MLP leads to a 20% drop in success rate. The SiT-based decoder (Ma et al. 2024) replaces the diffusion objective with a flow-matching one, yielding a lower average score than the DiT-based decoder.
@@
The impact of frame number. As shown in Fig.5, we analyze how varying the number of input frames affects different multi-frame modeling strategies. Given that tasks differ in their reliance on temporal information, we conduct a representative analysis on the SimplerEnv benchmark. The study compares CronusVLA 7B, CronusVLA 0.5B, and the Baseline model (post-trained basic single-frame VLA mode) with a naive multi-frame extension. Results show several conclusions: (1) Increasing the total frame number yields an initial increase followed by a subsequent decrease in average success rate. More frames do not consistently lead to better outcomes: CronusVLA 7B performs best with 7 frames, while the CronusVLA 0.5B performs best with 4, suggesting that a moderate amount of temporal information can enhance performance, whereas excessive temporal input may lead to performance degradation. (2) Compared with Baseline, CronusVLA maintains high inference speed across frame counts, avoiding excessive latency overhead. (3) CronusVLA demonstrates strong long-horizon efficiency, with inference speed of baseline models significantly degrading as the horizon extends, while our model remains largely unaffected.
@@ 5 Conclusion
We presented CronusVLA, a unified framework that efficiently extends single-frame VLA models to the multi-frame paradigm through a two-stage training process. By introducing feature chunking and multi-frame regularization, CronusVLA achieves fast inference, strong performance, and enhanced robustness. Additionally, we proposed SimplerEnv-OR, a benchmark for evaluating observational robustness under diverse disturbances. Extensive experiments across simulated and real-world settings validate the effectiveness and robustness of our approach. Limitations and Future Work are discussed in detail in Appendix.
Supplementary Materials
Appendix A SimplerEnv-OR
A.1 Benchmark Settings
Figure 6: SimplerEnv-OR, which includes different disturbance categories (Global, Local, and Discrete), with different disturbed-to-clean ratios (1:0 for Constant, 1:1 for Cyclic, 1:3 or 1:5 for Sparse), supporting spatial- and temporal-dimension evaluations.
SimplerEnv-OR (Observational Robustness) is designed to evaluate model robustness against observational disturbances across temporal and spatial dimensions. It builds upon the WidowX Robot setup from SimplerEnv (Li et al. 2024b), reusing the original Visual Matching configuration, which includes four subtasks: Stack Green Cube on Yellow Cube, Put Carrot on Plate, Put Spoon on Table, and Put Eggplant in Basket. These tasks are all pick-and-place tasks and exhibit stable success rates, making them well-suited for assessing model robustness. As in the original SimplerEnv, these tasks are evaluated in a real-to-sim setting using models trained on the BridgeV2 dataset (Walke et al. 2023). SimplerEnv-OR directly tests each model under various disturbance types without additional fine-tuning, and computes the average success rate across the four subtasks for each type. We report both the success rate and the predefined robustness score (R-Socre) over the corresponding disturbance types. The same random seeds are used to ensure that all baseline models encounter identical disturbances within the same episode and trial during evaluation. A visualization is shown in Fig.6.
Disturbance Spatial Category Temporal Category Ratio Supported Subtask Number Trials
Blurring Global Constant,Cyclic,Sparse 1:0, 1:1, 1:3 4 288
Jittering Global Constant,Cyclic,Sparse 1:0, 1:1, 1:3 4 288
Frame Dropping Global Cyclic,Sparse 1:1, 1:3, 1:5 4 288
Full Occlusion Global Cyclic,Sparse 1:1, 1:3, 1:5 4 288
Overexposing Local Cyclic,Sparse 1:1, 1:3, 1:5 4 288
Partial Occlusions Local Cyclic,Sparse 1:1, 1:3, 1:5 4 288
Gaussian Noise Discrete Constant,Cyclic,Sparse 1:0, 1:1, 1:3 4 288
Impulse Noise Discrete Constant,Cyclic,Sparse 1:0, 1:1, 1:3 4 288
@@ A.1 Benchmark Settings
Table 6: Task details of SimplerEnv-OR, including types of observational disturbances, corresponding spatial categories, temporal categories, supported ratios (i.e., Disturbed:Clean, frame number), subtask number, and evaluation trials number.
Disturbance Foundation Core Parameters Sampling Configurations
[…]
Jittering motion blur direction, kernel size direction $\in\{H,V,D\}$, size $\in\{15,25,50\}$
Frame Dropping frame replication temporal offset use past_image_list[-1]
Full Occlusion full mask (none) image $\leftarrow 0$
Overexposing exposure varying position, intensity, radius intensity $\in\{0.7,0.85,1.0\}$, radius based on image size
Partial Occlusions partial mask position, radius $\{1/4,2/4,3/4\}$width, radius based on image size
Gaussian Noise gaussian noise coordinates, std std $\in\{25,50,75\}$
Impulse Noise salt and pepper noise coordinates, amount, salt_vs_pepper amount $\in\{0.2,0.5,0.8\}$, s/p $\in\{0.0,0.5,1.0\}$
Table 7: Detailed configurations of disturbances, summary of visual disturbance types, foundation tools, core parameters, and sampling configurations.
A detailed summary of disturbances conducted in our SimplerEnv-OR is illustrated in Tab.6 and Tab.7. Notably, commonly used image augmentations during VLA models training, such as random resized cropping, adjustments to brightness, contrast, saturation, and hue, are beyond the scope of our benchmark, and the designed tasks in SimperEnv-OR cannot be directly solved through these simple data augmentation techniques.
A.2 Spatial Evaluations on SimplerEnv-OR
Figure 7: Spatial-dimension results on SimplerEnv-OR. The left panel presents the average success rates across all trials (under 8 major disturbance types, 3 spatial categories). The right panel shows the corresponding R-Score performances.
The spatial dimension assesses model robustness by injecting various types of disturbances into clean observations, simulating realistic interference encountered by cameras in practical deployments. Observation disturbances are typically categorized into three categories: Global, Local, and Discrete. Global disturbances affect the entire observation, resulting in complete information loss (e.g., Full Occlusion) or uniform degradation (e.g., Blurring). Local disturbances are confined to specific regions of the observation, leaving the remaining areas unaffected, such as Partial Occlusions. Discrete disturbances refer to randomly distributed, non-contiguous noise patterns where only a subset of pixels within a region are affected, characterizing sparse and irregular corruption.
As illustrated in Fig. 7, the left and right panels show the average success rates and R-Scores, respectively, across Global, Local, and Discrete in SimplerEnv-OR. Our CronusVLA achieves the highest success rate under nearly all disturbance types. It also demonstrates superior robustness in R-Score, particularly under Full Occlusion, Partial Occlusion, Jittering, and Blurring, while maintaining competitive performance across other types. This robustness can be attributed to the temporal modeling of our multi-frame method. SpatialVLA shows strong robustness to Frame Dropping and Blurring but is highly sensitive to other disturbances, especially Full Occlusion, likely due to its reliance on zero-shot depth estimation. Although RoboVLMs and TraceVLA perform less effectively than SpatialVLA in the original SimplerEnv, they exhibit enhanced robustness on SimplerEnv-OR. CogACT achieves the best R-Score performance on several disturbance types but falls significantly behind CronusVLA in Full Occlusion, Partial Occlusion, Blurring and Jittering. We also test both the JAX and the Lerobot version of $\pi_{0}$.
Figure 8: Visualizations of spatial-dimension testing on SimplerEnv-OR. Qualitative comparisons under spatial disturbances such as Cyclic Global Full Occlusion (upper) and Cyclic Local Partial Occlusion (bottom).
Method Constant (1:0) Cyclic (1:1) Sparse (1:3) Sparse (1:5)
Global-p Discrete Avg. Global Discrete Local Avg. Global Discrete Local Avg. Global-p Local Avg.
$\pi_{0}$(jax) 7.3 10.9 9.1 8.3 11.5 2.6 7.7 7.6 10.4 3.7 7.3 14.1 11.5 12.8
lerobot-$\pi_{0}$ 27.1 32.8 30.0 21.6 31.8 6.8 20.4 22.9 33.3 8.3 21.9 39.1 28.1 33.6
TraceVLA 11.5 21.4 16.4 16.4 22.4 14.1 17.3 20.1 24.5 21.9 21.6 12.0 18.2 15.1
SpatialVLA 22.9 18.2 20.6 25.8 21.9 14.6 22.0 28.9 31.8 26.0 28.9 26.0 28.1 27.1
RoboVLMs 16.7 24.0 20.3 20.3 31.8 25.5 24.5 26.6 42.7 38.5 33.6 29.7 42.7 36.2
CogACT 17.2 41.7 29.4 28.4 52.1 37.0 36.5 41.2 51.0 43.8 44.3 43.2 52.6 47.9
[…]
Table 8: Spatial-dimension success rates under different ratios. For each disturbed-to-clean ratio (from 1:0 to 1:5), we report all success rates. Global-p indicates that a part of the Global disturbances are included.
Method Constant (1:0) Cyclic (1:1) Sparse (1:3) Sparse (1:5)
Global-p Discrete Avg. Global Discrete Local Avg. Global Discrete Local Avg. Global-p Local Avg.
$\pi_{0}$(jax) 34.9 52.3 43.6 39.9 54.8 12.5 36.8 36.1 49.8 17.4 34.9 67.3 54.8 61.1
lerobot-$\pi_{0}$ 54.2 65.6 59.9 43.2 63.5 13.5 40.9 45.8 66.7 16.7 43.8 78.1 56.2 67.2
TraceVLA 41.4 77.1 59.2 59.2 80.9 50.8 62.5 72.4 88.4 79.0 78.0 43.3 65.8 54.5
SpatialVLA 50.0 39.8 44.9 56.3 47.8 31.8 48.1 63.1 69.4 56.9 63.1 56.9 61.4 59.1
RoboVLMs 39.0 56.1 47.6 47.6 74.4 59.8 57.3 62.2 100.0 90.3 78.7 69.5 100.0 84.8
CogACT 31.1 75.5 53.3 51.4 94.4 67.0 66.1 74.5 92.5 79.3 80.2 78.3 95.3 86.8
[…]
Table 9: Spatial-dimension R-Score under different ratios. For each disturbed-to-clean ratio (from 1:0 to 1:5), we report the R-Score corresponding to different spatial-level categories. Global-p indicates that a part of the Global disturbances are included.
Fig. 8 presents qualitative results. In the upper, Stack Green Cube on Yellow Cube task under Global (Full Occlusion) and Cyclic disturbances, SpatialVLA frequently outputs out-of-domain actions and feature representation, leading to task failures. CogACT, also relying solely on the current frame, produces contextually irrelevant actions when occlusion occurs. In contrast, CronusVLA, equipped with multi-frame modeling, demonstrates superior robustness under such conditions. In the bottom, Put Carrot on Plate task under Local (Partial Occlusion) and Cyclic disturbances, RoboVLMs, although leveraging multi-frame modeling, depend heavily on accurate temporal information. When this information is compromised, the model avoids out-of-distribution actions but repeats ineffective attempts. CogACT misinterprets the scene, interacting with irrelevant objects (e.g., the table), resulting in task failure. CronusVLA maintains robust performance under Partial Occlusion by effectively integrating temporal context.
We report more detailed experimental results in Tab.8 and Tab.9. All values represent the average performance across all disturbance types within the respective spatial category. Our model demonstrates clear advantages over others in both Global and Local evaluations, and exhibits strong robustness in the Discrete setting.
A.3 Temporal Evaluations on SimplerEnv-OR
The temporal dimension assesses model robustness by injecting disturbances at varying intervals, simulating different disturbance ratios in otherwise clean observations. This design captures realistic disturbance patterns encountered by cameras in practical deployment scenarios, ranging from persistent to intermittent interference caused by external factors. To enable a meaningful qualitative analysis, we evaluate three representative settings: Constant (1:0), Cyclic (1:1), and Sparse (1:3, 1:5).
Firstly, to better illustrate the impact of different disturbance ratios on model performance, we present detailed scores under Full Occlusion in Tab.10, along with results from the original SimplerEnv across all four subtasks. As shown in this table, success rates generally improve as the disturbance ratio decreases. Under Cyclic disturbances, most models exhibit near-zero success rates, as alternating normal and fully occluded frames severely disrupt single-frame decision logic, often resulting in out-of-distribution actions. In contrast, our CronusVLA, equipped with temporal modeling, effectively leverages past observations to infer correct decisions, demonstrating superior robustness.
Method Cyclic (1:1) Sparse (1:3)
Spoon Carrot Blocks Eggplant Avg. R-Score Spoon Carrot Blocks Eggplant Avg. R-Score
SpatialVLA 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0
RoboVLMs 0.0 0.0 0.0 0.0 0.0 0.0 41.7 12.5 0.0 4.2 14.6 30.5
CogACT 8.3 0.0 0.0 8.3 4.2 5.4 25.0 25.0 4.2 54.2 27.1 44.2
[…]
Sparse (1:5) WR-VM (SimplerEnv)
Spoon Carrot Blocks Eggplant Avg. R-Score Spoon Carrot Blocks Eggplant Avg. -
SpatialVLA 0.0 8.3 0.0 29.2 9.4 14.3 20.8 37.5 41.7 83.3 45.8
RoboVLMs 33.3 20.8 0.0 29.2 20.8 39.3 50.0 37.5 0.0 83.3 42.7
CogACT 37.5 29.2 0.0 75.0 35.4 50.8 75.0 50.0 16.7 79.2 55.2
[…]
Table 10: Detailed experimental results under Full Occlusion across three disturbance ratios: 1:1, 1:3, 1:5. Origin results of SimplerEnv WR-VM setting is also reported. Success rates (%) are reported. “R-Score” indicates the robustness score.
We present qualitative results in Fig.9, including a Constant disturbance under Jittering and a Sparse disturbance under Impulse Noise. In the upper example, Constant disturbances significantly impact model performance: RoboVLMs exhibit incorrect position perception due to insufficient accurate historical information, which may be caused by its LSTM-based design. CogACT fails to maintain object grasping under disturbance. In the bottom example, SpatialVLA suffers from wrong depth estimations, resulting in failed grasps at incorrect heights, while RoboVLMs show execution hesitation, reducing success rates. In contrast, CronusVLA demonstrates higher success rates under extreme conditions (Constant) and remains largely unaffected under mild disturbances (Sparse).
Fig. 10 illustrates the impact of varying types and ratios of disturbances on different models. Some disturbance types exhibit increased success rates at higher disturbance ratios. We attribute this to two main factors: (1) inherent randomness in both the models and disturbances, as well as task-specific variance, which makes performance more stable when averaged over more trials; (2) disturbances during critical moments may disproportionately affect model performance, and higher ratios do not necessarily increase interference at those moments. SimplerEnv-OR currently adopts fixed-frequency disturbances, which could be improved by incorporating targeted perturbations on keyframes for more accurate robustness evaluation.
Figure 9: Visualizations of temporal-dimension testing on SimplerEnv-OR. Qualitative comparisons under spatial disturbances such as Constant Global Jittering (upper) and Sparse Discrete Impulse Noise (bottom).
Figure 10: Impact of disturbance types and ratios on models performance. Both the success rate and R-Score are reported.
A.4 More Discussions
Our design features three key properties:
(1) Plug-and-Play and Transferable. We focus on evaluating the robustness of observations, which are typically obtained via environment rendering. To ensure decoupling from simulation environments, we introduce disturbances directly on rendered images (as NumPy arrays), without modifying the simulator itself. This approach is compatible with any simulation or standard benchmark.
(2) Extensibility. The types and intensities of disturbances are fully customizable without additional configuration, allowing seamless integration of various image processing techniques.
(3) Reproducibility. We ensure consistency across testing trials by deterministically mapping each trials id to a random seed using its hash value, guaranteeing identical disturbances for different models during evaluation.
However, our approach still has several limitations, such as the incomplete coverage of disturbance types, which we leave for future work. Expanding the number of baselines, supporting more simulation environments, and incorporating more accurate and comprehensive evaluation protocols are also our future directions.
We hope to inspire the community to consider robustness evaluation in Vision-Language-Action (VLA) models, which is crucial for VLA real-world deployment and future applications.
Temporal Dimension WidowX Robot
Constant (1:0) Cyclic (1:1) Sparse (1:3) VM
Methods R-Score SR R-Score SR R-Score SR - SR
$\pi_{0}$(jax) (Black et al. 2024) 43.5 9.1 36.8 7.7 34.9 7.3 - 20.9
lerobot-$\pi_{0}$ (Cadene et al. 2024) 59.9 30 40.9 20.4 43.8 21.9 - 50.0
TraceVLA (Zheng et al. 2024) 59.2 16.4 62.5 17.3 78.0 21.6 - 27.7
RoboVLMs (Li et al. 2024c) 47.6 20.3 57.3 24.5 78.7 33.6 - 42.7
SpatialVLA (Qu et al. 2025) 44.9 20.6 48.0 22.0 63.1 28.9 - 45.8
CogACT (Li et al. 2024a) 53.3 29.4 66.1 36.5 80.2 44.3 - 55.2
[…]
Spatial Dimension Total Avg.
Global Local Discrete
Methods R-Score SR R-Score SR R-Score SR R-Score SR
$\pi_{0}$(jax) (Black et al. 2024) 42.6 8.9 28.2 5.9 52.1 10.9 41.1 8.6
lerobot-$\pi_{0}$ (Cadene et al. 2024) 51.7 25.9 28.8 14.4 65.3 32.6 49.4 24.7
TraceVLA (Zheng et al. 2024) 58.1 16.1 65.3 18.1 81.9 22.7 65.8 18.2
RoboVLMs (Li et al. 2024c) 54.7 23.4 83.3 35.6 76.8 32.8 67.4 28.8
SpatialVLA (Qu et al. 2025) 57.6 26.4 50.0 22.9 52.4 24.0 54.4 24.9
CogACT (Li et al. 2024a) 60.2 33.2 80.5 44.4 87.4 48.3 72.1 39.8
[…]
Table 11: Additional results on SimplerEnv-OR. Top: Temporal dimension with different frequencies. Bottom: Spatial dimension with different patterns. R-Score indicates the predefined robustness score. SR denotes success rate. $\pi_{0}$ models on different frameworks (JAX or Torch) are included.
Appendix B Case Study for Long-horizon Tasks
B.1 Simulation Experiments
Our model achieves a considerable performance on the Put in Drawer task of SimplerEnv (Li et al. 2024b) benchmark, reaching 64.8% and 65.1% in Visual Matching and Variant Aggregation, compared to baseline methods including OpenVLA (Kim et al. 2024), SpatialVLA (Qu et al. 2025), RoboVLMs (Li et al. 2024c) and TraceVLA (Zheng et al. 2024), which achieve only 0%–30% success rates. As shown in Fig.11, we present a case study in the SimplerEnv setting with a Google Robot. In (A) and (B), OpenVLA struggles to fully open the drawer, blocking progress at the first step. In contrast, CronusVLA consistently completes all steps of the long-horizon task. In (C) and (D), although RoboVLMs can successfully open the top drawer initially, they fail to maintain the new execution goal, resulting in inaccurate actions and repeated collisions. This may stem from inadequate modeling capacity to adapt execution goals throughout long-horizon tasks. CronusVLA, by contrast, executes all steps smoothly and reliably.
B.2 Real-world Experiments